# Antibunched emission from femtosecond laser-written nano-ablations on gallium nitride

### v1

@@ I Introduction
In the last decade, quantum emitters (QEs) in gallium nitride (GaN) have been investigated as a practical source of quantum light at room temperature[1]. These QEs display low multi-photon emission probability[2], comparatively high Debye-Waller factor[3], high continuous wave (CW) photon detection rate[4] and telecom range emission[5]. Recently, resonant optical excitation[6] and optically detected magnetic resonance have been observed in GaN QEs[3, 7], pointing to potential applications in quantum sensing and quantum networks. However, the reliable fabrication of QEs in GaN remains a challenge. This greatly hinders high-yield coupling to photonic structures that manipulate the local density of optical states and enhance photon collection[8].
Previous approaches to creating these QEs, such as ion implantation and doping during growth[9, 10, 11], create QEs with random positions and residual lattice damage[12, 13] which one may expect to degrade the spin and optical coherence properties. Recently, femtosecond laser-writing fabrication has attracted intense attention for its precise and highly controlled creation of QEs in semiconductors[14], such as diamond[15], silicon carbide [16], hexagonal boron nitride[17, 18], and aluminum nitride[19]. Moreover, femtosecond laser-writing is also used as a promising method to fabricate 3D photonic circuits from nanoscale to microscale[12, 20, 13]. Direct laser-writing offers a maskless fabrication process below the optical diffraction limit. Under optimal fabrication conditions, laser-written waveguide-integrated QEs in diamond feature spin coherence properties comparable to their native counterparts in the host material[21, 22].
In this paper, we report femtosecond laser-writing and subsequent annealing to fabricate an array of emitters in GaN. Laser-written nano-ablations in GaN are mapped by atomic force microscopy (AFM) and confocal photoluminescence (PL) mapping. Compared to previously studies which reported broadband PL emission[23, 24], the laser-written structures in this paper exhibit a sharp emission peak in their spectrum at room temperature, antibunching in their photon emission correlation spectra (PECS), shelving behavior in time-resolved PL (TRPL) measurements, and few saturation PL rates under CW excitation. These features are consistent with the creation of a small number of QEs in a sub-diffraction-limited spot. Our study provides a promising route to scalable, integrated, on-chip quantum technologies based on GaN via femtosecond laser writing.
@@ II.2 Laser-written bright antibunched emitters
To understand these emitters’ bunching behavior and energy structure[25], a TRPL spectrum is recorded under two duration laser pulse excitation with , as shown in Fig. 1(i). The spacing between each double-pulse train is to allow the ground state population to reset. From the shelving behavior during the laser pulse excitation, we deduce that at least three-energy-levels are present in these emitters, which includes the ground state (GS), excited state (ES), and shelving state (SS). In Figs. 1(j), the excited power-dependent PL intensity is fitted by
@@ Acknowledgements.
The authors acknowledge financial support provided by EPSRC via Grant No. EP/T017813/1 and EP/03982X/1 and the European Union’s H2020 Marie Curie ITN project LasIonDef (GA No. 956387) and Sêr Cymru National Research Network in Advanced Engineering and Materials. IFN-CNR is thankful for support from the projects QuantDia (FISR2019-05178) and PNRR PE0000023 NQSTI funded by MUR (Ministero dell’Università e della Ricerca). Device processing was carried out in the cleanroom of the ERDF-funded Institute for Compound Semiconductors (ICS) at Cardiff University. We wish to thank ICS staff, Miguel Alvarez Perez and Katie Eggleton for technical support.Author Contributions
@@
Conceptualization: YG, AJB, SME; Methodology: YG; Software: YG; Data curation: YG; Formal Analysis: YG; Investigation: YG, GC; Resources: All; Writing - Original Draft: YG; Writing - Review & Editing: All; Supervision: SME, JPH, AJB; Funding Acquisition: JPH, SME, AJB.
@@
Supplemental Materials: Antibunched emission from femtosecond laser-written nano-ablations on gallium nitride
I Confocal ODMR setup
@@ II AFM measurement
The AFM measurement was conducted in commercial Brucker AFM microscopes. The scanning tip is the silicon tip on the nitride lever. The scan parameters are of scanning rate with 256 samples and of scan angle. The AFM image was analysed and plotted by the Brucker commercial software (Nanoscope Analysis 3.0).

### v2

# Femtosecond laser-written nano-ablations containing bright antibunched emitters on gallium nitride

@@ I Introduction
In the last decade, quantum emitters (QEs) in gallium nitride (GaN) have been investigated as a practical source of quantum light at room temperature[1]. These QEs display low multi-photon emission probability[2], comparatively high Debye-Waller factor[3], high continuous wave (CW) photon detection rate[4] and telecom range emission[5]. Recently, resonant optical excitation[6] and optically detected magnetic resonance have been observed in GaN QEs[3, 7], pointing to potential applications in quantum sensing and quantum networks. However, the reliable fabrication of QEs in GaN remains a challenge. This greatly hinders high-yield coupling to photonic structures that manipulate the local density of optical states and enhance photon collection[8], which is crucial for the realization of scalable integrated quantum photonics.
Previous approaches to creating these QEs, such as ion implantation and doping during growth[9, 10, 11, 2, 3], create QEs with random positions and residual lattice damage[12, 13] which one may expect to degrade the spin and optical coherence properties. Recently, femtosecond laser-writing fabrication has attracted intense attention for its precise and highly controlled creation of QEs in semiconductors, such as diamond[14, 15, 16], silicon carbide [17], hexagonal boron nitride[18, 19], and aluminum nitride[20]. Moreover, femtosecond laser-writing is also used as a promising method to fabricate 3D photonic circuits from nanoscale to microscale[12, 21, 13]. Direct laser-writing offers a maskless fabrication process below the optical diffraction limit. Under optimal fabrication conditions, laser-written waveguide-integrated QEs in diamond feature spin coherence properties comparable to their native counterparts in the host material[22, 23].
In this paper, we report femtosecond laser-writing and subsequent annealing to fabricate an array of emitters in GaN. Because the bright spots observed in the PL scan are in a regular square array, we can be certain that the laser writing induces them, and not intrinsic to the semiconductor. As we will show, bright emission is correlated in position with the nano-ablated holes on the sample surface. Laser-written nano-ablations in GaN are mapped by atomic force microscopy (AFM) and confocal photoluminescence (PL) mapping. Compared to previously studies which reported broadband PL emission[24, 25], the laser-written structures in this paper exhibit a sharp emission peak in their spectrum at room temperature, antibunching in their photon emission correlation spectra (PECS), shelving behavior in time-resolved PL (TRPL) measurements, and few saturation PL rates under CW excitation. These features are consistent with the creation of a small number of QEs in a sub-diffraction-limited spot. Our study provides a promising route to scalable, integrated, on-chip quantum technologies based on GaN via femtosecond laser writing.
@@ II.2 Laser-written bright antibunched emitters
To understand these emitters’ bunching behavior and energy structure[26], a TRPL spectrum is recorded under two duration laser pulse excitation with a space, as shown in Fig. 1(i). The spacing between each double-pulse train is to allow the ground state population to reset. From the shelving behavior during the laser pulse excitation, we deduce that at least three energy levels are present in these emitters, which include the ground state (GS), excited state (ES), and shelving state (SS), consistent with the previous study[27]. In Figs. 1(j), the excited power-dependent PL intensity is fitted by
@@ Acknowledgements.
The authors acknowledge financial support provided by EPSRC via Grant No. EP/T017813/1 and EP/03982X/1 and the European Union’s H2020 Marie Curie ITN project LasIonDef (GA No. 956387) and Sêr Cymru National Research Network in Advanced Engineering and Materials. IFN-CNR is thankful for support from the projects QuantDia (FISR2019-05178) and PNRR PE0000023 NQSTI funded by MUR (Ministero dell’Università e della Ricerca). Device processing was carried out in the cleanroom of the ERDF-funded Institute for Compound Semiconductors (ICS) at Cardiff University. We wish to thank ICS staff and Miguel Alvarez Perez for technical support.Author Contributions
@@
YG: Conceptualization, Methodology, Software, Data curation, Formal Analysis, Investigation, Resources, Writing - Original Draft, Writing - Review & Editing. GC: Investigation, Resources, Writing - Review & Editing. VB: Investigation, Resources, Writing - Review & Editing. RY: Investigation, Resources, Writing - Review & Editing. KE: Investigation, Writing - Review & Editing. JPH: Conceptualization, Resources, Writing - Review & Editing, Supervision, Funding Acquisition. AJB: Conceptualization, Resources, Writing - Review & Editing, Supervision, Funding Acquisition. SME: Conceptualization, Resources, Writing - Review & Editing, Supervision, Funding Acquisition.
@@
Conflict of Interest
The authors declare no conflicts of interest.
Supplemental Materials: Femtosecond laser-written nano-ablations containing bright antibunched emitters on gallium nitride
I Confocal setup
@@ II AFM measurement
The AFM measurement was conducted in commercial Bruker AFM microscopes. The scanning tip is the silicon tip on the nitride lever. The scan parameters are of scanning rate with 256 samples and of scan angle. The AFM image was analysed and plotted by the Brucker commercial software (Nanoscope Analysis 3.0).